OpenStudy (anonymous):

WILL SOMEONE PLEASE HELP? Please show all work:: Perform the indicated operations:: Let f(x) = x2 - 7x - 9. Find f(a + h) - f(a) and simplify.

OpenStudy (anonymous):

f(a)=a^2-7a-9 f(a+h)=a^2+2ah+h^2-7a-7h-9 (a+h)^2=a^2+2ah+h^2 just so you know 2ah+h^2-7h=f(a+h)-f(a)

OpenStudy (anonymous):

so the final answer is 2ah+g^2-7h=f(a+h)-f(a) ?....

OpenStudy (anonymous):

h^2 not g^2 but yea.

OpenStudy (anonymous):

oops, typo- thank you!

OpenStudy (anonymous):

but if you simplify its h(2a+h-7)

OpenStudy (anonymous):

ok, great- thanks!